# on CLIP via Concept Confusion

### v1

@@ Abstract
Backdoor attacks pose a significant threat to deep learning models, enabling adversaries to embed hidden triggers that manipulate the behavior of the model during inference. Traditional backdoor attacks typically rely on inserting explicit triggers (e.g., external patches, or perturbations) into input data, but they often struggle to evade existing defense mechanisms. To address this limitation, we investigate backdoor attacks through the lens of the reasoning process in deep learning systems, drawing insights from interpretable AI. We conceptualize backdoor activation as the manipulation of learned concepts within the model’s latent representations. Thus, existing attacks can be seen as implicit manipulations of these activated concepts during inference. This raises interesting questions: why not manipulate the concepts explicitly? This idea leads to our novel backdoor attack framework, Concept Confusion Attack ($C^{2}$ttack), which leverages internal concepts in the model’s reasoning as ”triggers” without introducing explicit external modifications. By avoiding the use of real triggers and directly activating or deactivating specific concepts in latent spaces, our approach enhances stealth, making detection by existing defenses significantly harder. Using CLIP as a case study, experimental results demonstrate the effectiveness of $C^{2}$ttack, achieving high attack success rates while maintaining robustness against advanced defenses.
@@ 1 Introduction
As a powerful multimodal model, CLIP [42] excels in understanding and categorizing images based on text descriptions through contrastive learning. By aligning images and textual descriptions in a shared embedding space, CLIP demonstrates impressive generalization across various vision-language tasks and domains [46]. Despite its success, CLIP presents significant security vulnerabilities, particularly in the form of backdoor attacks [9]. In a backdoor attack, adversaries embed hidden triggers during training that allow them to manipulate the model behavior when specific inputs or triggers are present while maintaining normal performance on clean data. Traditional backdoor attacks often involve inserting external patches into images as triggers [30, 8, 36]. More advanced approaches, such as BadCLIP [2] and SSBA [27], generate learnable and imperceptible additive noise as backdoor triggers. Some attacks, such as physical backdoor attacks [45, 1], use naturally occurring or artificially introduced attributes as triggers. These can include specific object appearances (e.g., cars painted in green), or physical markers embedded in real-world scenes. However, such approaches often require the artificial attribute to be visually dominant, limiting their applicability in complex backgrounds such as forests or lakes. Furthermore, these methods still necessitate explicit trigger injection and remain susceptible to existing defense mechanisms [48]. This raises a natural question: Can we exploit backdoor attacks without explicitly injecting external triggers while also evading detection by current defense mechanisms?
To answer the question, we first fundamentally understand the inner workings of backdoor attacks through the lens of reasoning mechanisms in deep learning systems. The Hopfieldian view in cognitive neuroscience [19, 3] provides valuable insights by emphasizing distributed representations and dynamic processes across neural populations. It explains reasoning as transformations within “representation spaces” in response to stimuli, offering a conceptual and semantic perspective on cognitive systems.
@@
Figure 1: Illustration of traditional backdoor attacks, physical attacks, and our $C^{2}$ttack. Traditional backdoor attacks inject external triggers, either visible or imperceptible, into images to manipulate model predictions. Physical attacks [45] rely on distinct physical objects or explicit visual attributes as triggers, making them externally visible in real-world settings. In contrast, our $C^{2}$ttack introduces no external triggers. Instead, it exploits human-understandable concepts that the model naturally relies on for decision-making, designating them as internal triggers. This approach makes $C^{2}$ttack more stealthy and resilient against conventional defenses.
These insights have inspired significant research in understanding deep learning models through their “distributed latent space representations.” Explainable AI methods aim to uncover what neural networks learn during training by analyzing representations such as visual features and human-interpretable concepts. For example, concept extraction methods provide insight into the features or human-understandable concepts that neurons or model layers respond to [51, 22, 12, 16]. Similar patterns have been observed across domains, including Natural Language Processing [41, 38].
Viewing backdoor-triggered decisions through this lens is intuitive. Backdoor triggers, like external stimuli, induce shifts in a model’s decision-making trajectory by altering activated representations without modifying the underlying system. Existing backdoor attacks implicitly manipulate activated features during inference by introducing external triggers. This raises a key question: Why not manipulate the concepts explicitly? Could the inherent concepts in a model’s reasoning serve as an “internal trigger” for the backdoor?
Motivated by this, we propose the Concept Confusion Attack ($C^{2}$ttack), which explicitly manipulates representations within the model’s latent space. By leveraging internal concepts, our approach eliminates the need for external triggers, enhances stealth, and evades traditional defenses focused on input anomalies. See Figure 1 for some backdoor examples.
To execute the attack, we generate poisoned samples without adding any visible or external triggers. Our method begins by extracting human-understandable concepts that the model relies on for decision making, such as “trees” or “cars”, and designating these as internal triggers. The images associated with these concepts are relabeled, whereas the visual content of the images remains unchanged. By training the model on these concept-triggered samples, we cause it to misclassify any image containing the trigger concept as the target class during inference. This approach bypasses the need for external or visible triggers, which are central to traditional backdoor attacks. Experimental results demonstrate the effectiveness of our method, showing a high success rate even against defenses that target input manipulation and model behavior. Our contributions are as follows.
  1. 1.
    We establish a connection between the decision-making process under backdoor attacks and reasoning process via explainable AI.
  2. 2.
    Based on these connections, we leverage the latent space representations of neural networks to deepen our understanding and improve backdoor attacks. We first examine how existing backdoor attacks function by manipulating the “implicit concepts” within the model’s latent representation. Building on this, we introduce a new framework, Concept Confusion Attack ($C^{2}$ttack), which avoids the need for physical triggers by directly manipulating “explicit concepts” within the model’s representation spaces.
  3. 3.
    Comprehensive experiments on several datasets and various defense methods demonstrate that our $C^{2}$ttack is harder to detect than traditional attacks that rely on input anomalies, achieving state-of-the-art performance compared to baseline methods.
@@ 2 Related Works
Backdoor Attack against CLIP. Recent research has expanded the scope of backdoor attacks to multimodal domains, illustrating their adaptability across diverse architectures. Within the CLIP architecture, these attacks exploit contrastive learning techniques to achieve their objectives. In particular, an early data poisoning attack on CLIP [8] aimed to force specific inputs to be misclassified with a targeted label. Similarly, Yang et al. [47] proposed a method that adjusts encoders to increase the cosine similarity between image and text embeddings, leading to misclassification in image-text retrieval tasks. Furthermore, BadEncoder [21] and BadCLIP [31] introduce backdoors by enhancing the similarity between poisoned image embeddings and target image embeddings. Another version of BadCLIP [2] injects a learnable trigger into both the image and text encoders during the prompt learning phase. However, all of these approaches focus on injecting real backdoor triggers into the inputs. Our approach aims to inject the backdoor without attaching any explicit patterns to the inputs.
Concept-based Explanations. Recent studies on the internal workings of deep neural network models have revealed fascinating properties in their learned representations. The linear representation hypothesis in explainable AI suggests that linear combinations of neurons can encode meaningful and insightful concepts, while neural networks frequently represent high-level features in linear directions in the activation space. Most current approaches focus on representation-level analysis and consider how these representations connect to concepts learned during pre-training [6, 44]. Other works explore concept localization and representation in neural networks [22, 26] and linear classifier probing to uncover input properties [5], fact localization and model editing [37, 50, 10, 11], concept erasure [43, 15], and corrective analysis [7]. These findings align with RepE [52], which emphasized the near-linear nature of representations in transformer-based architectures [41]. In parallel, Concept Bottleneck Models (CBMs) have introduced concept bottleneck layers into deep neural networks, improving model generalization and interpretability by learning specific, meaningful concepts [23]. Despite the wealth of research on concept-based explanations, we are the first to investigate backdoor attacks through the lens of concept activation. We conceptualize the decision process triggered by backdoors as the movement between representations of learned concepts. By focusing on manipulating these learned concepts, our approach opens new avenues for understanding and exploiting a model’s internal representations for backdoor attacks.
@@
Table 1: Top-5 concepts extracted from single attention heads of CLIP-ViT-L/14 during clean training and backdoor training (with BadNet [17]) on CIFAR-10, where L represents transformer layers and H denotes attention heads. Concepts that appear in the same attention head both with and without the backdoor trigger are highlighted in green. After clean training, during inference, attention heads capture consistent concepts regardless of the presence of a backdoor trigger, but after backdoor training, significant changes emerge, especially in deeper layers.
@@
3 Preliminaries
3.1 Problem Definition
CLIP [42] is a multimodal model with vision and language encoders trained on large-scale image-caption data. This paper focuses on its vision encoder for image classification, though our attack framework can be extended to other vision-language models.
CLIP-based Image Classification. Suppose $D=\{(x_{1},y_{1}),\cdots,(x_{N},y_{N})\}$ is an image dataset consisting of $N$ training samples, where $x_{i}\in\mathcal{X}$ is the $i$-th image and $y_{i}\in\mathcal{Y}$ is its corresponding label. Let $f:\mathcal{X}\rightarrow\mathcal{E}$ denote a CLIP vision encoder that maps any image from the sample space $\mathcal{X}$ to the embedding space $\mathcal{E}$. To perform classification with the CLIP vision encoder $f$, one will first construct an image classification model based on the encoder $f$ as $g:=h(f(\cdot)):\mathcal{X}\rightarrow\mathcal{Y}$, where $h:\mathcal{E}\rightarrow\mathcal{Y}$ is a classification head mapping encoded representations to their predicted labels. Then, the overall classification model $g$ will be fine-tuned (i.e., both $f$ and $h$ will be fine-tuned) in the training set $D$ by minimizing the objective function $\mathcal{L}(f,h,D):=\frac{1}{N}\sum_{i=1}^{N}\ell(h(f(x_{i})),y_{i})$, where $\ell:\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R}^{+}$ is a loss function.
3.2 Threat Model
We investigate targeted backdoor attacks on CLIP-based classification models.
Adversary’s Goal. The adversary aims to train a backdoored model without explicitly injecting external trigger patterns into input images. The model behaves normally on clean images, but misclassifies poisoned images, which inherently contain strong semantic concepts that influence the model’s decision-making. Unlike existing backdoor attacks [17, 9, 39], which rely on explicit trigger injection, our approach constructs poisoned data without modifying the original images. Following the standard threat model [17], we assume that the adversary has full control over the training process and the construction of poisoned data.
Formal Definition. Let $D^{(p)}=\{(x^{(p)}_{1},y_{\mathrm{target}}),\dots,(x^{(p)}_{M},y_{\mathrm{target}})\}$ be a poisoned dataset containing $M$ samples, where each image $x^{(p)}_{i}$ inherently exhibits specific concepts $P$, and its label is assigned to a pre-defined target label $y_{\mathrm{target}}\in\mathcal{Y}$ [^1] Details on constructing the poisoned dataset are provided in Sec. 4.2. The adversary injects $D^{(p)}$ into the clean dataset $D$, forming a backdoored training set $\hat{D}:=D^{(p)}\cup D$. Training a CLIP-based classification model $g:=h(f(\cdot))$ on $\hat{D}$, following the procedure in Sec. 3.1, results in a backdoored model. For any poisoned image $x^{(p)}$ containing specific concepts $P$, the backdoored model predicts the target label i.e., $g^{*}(x^{(p)})=y_{\mathrm{target}}$. Meanwhile, for any clean image $x$ without $P$, the model maintains normal behavior, predicting labels consistent with clean training.
@@ 4 Concept Confusion Framework
As discussed in the Introduction section, recent advances in explainable AI have revealed that latent representations derived from activations across network layers can be decoded into meaningful and human-interpretable features/concepts. Building upon this understanding, we hypothesize that, unlike clean training, the attack-triggering process in backdoor attacks activates some distinct representations or learned concepts during the inference phase compared to clean samples. In Sec. 4.1, we conduct preliminary experiments and analyze the changes of activated concepts in the latent presentation of model layers, comparing the behavior of models between the clean and backdoor training processes. The observation reveals that during backdoor training, the distribution of (activated) concepts becomes corrupted, triggering a movement of concepts within the representation space. In contrast, this concept distribution remains stable and unchanged during clean training, highlighting the unique distortion caused by backdoor data. These results highlight that the decision process elicited by the backdoor-activation is a manipulation of activated concepts. Based on these understandings, in Sec. 4.2, we propose our Concept Confusion Attack.
4.1 Backdoor Attack Observations
To verify our previous conjecture, we conduct experiments to visualize and illustrate their underlying mechanisms. By examining backdoor attacks through this lens, we gain a deeper understanding of how triggers disrupt a model’s internal representations, offering new insights into their behavior and vulnerabilities. Specifically, we conceptualize the decision process triggered by backdoor attacks as a shift between activated representation (i.e., learned concepts). In this view, current backdoor methods implicitly manipulate the activated concepts during inference.
As a case study, we analyze the impact of BadNet [17], a widely used backdoor attack, on concept perception in CLIP-based models, with results shown in Tab. 2. To inject a backdoor trigger into a given image, BadNet will modify a small part of the pixels in the image to white/black pixels as the trigger pattern. In addition, the concepts adopted are label names from the Broden dataset [4]. We first finetune two classifiers built upon CLIP-ViT-L/14 [42] on the clean and backdoored (backdoored via BadNet) CIFAR-10 [24] datasets respectively and then leverage TextSpan [14], an algorithm designed for CLIP models to decompose concepts perceived by different attention heads in CLIP, to analyze how BadNet would affect concepts perceived by CLIPs.
The results shown in Tab. 2 reveal a significant impact of BadNet on the CLIP encoder’s internal representations. After clean training, the concepts captured by the attention heads remain largely consistent for inputs with or without the backdoor trigger. However, after backdoor training, comparing samples with and without backdoor triggers, dramatic changes occur, particularly in deeper layers: the 15th head in the 20th layer and the 1st head in the 23rd layer capture entirely different concepts, while the 5th head in the 22nd layer retains only the “Back_pillow” concept. This concentration of changes in later layers suggests that the backdoor attack primarily influences high-level abstractions and decision-making processes. The profound alterations in these attention heads indicate that the backdoor trigger induces substantial deviations in concept activation, likely explaining the attack’s high success rate while maintaining clean accuracy. These findings illuminate the mechanism by which concepts are altered within CLIP attention heads under backdoor attacks, providing insight into how such attacks manipulate model behavior. This confirms our hypothesis that backdoor-activation induces movement between representation spaces, while clean training maintains concept stability. Building on this understanding, we propose a novel method that explicitly manipulates concepts. Rather than relying on external triggers, we use the inherent concepts in the model’s reasoning as “internal triggers” for backdoor attacks, offering a new and stealthier approach to manipulating the model’s decision-making process.

### v2

# Backdooring CLIP through Concept Confusion

@@ Abstract
Backdoor attacks pose a serious threat to deep learning models by allowing adversaries to implant hidden behaviors that remain dormant on clean inputs but are maliciously triggered at inference. Existing backdoor attack methods typically rely on explicit triggers such as image patches or pixel perturbations, which makes them easier to detect and limits their applicability in complex settings. To address this limitation, we take a different perspective by analyzing backdoor attacks through the lens of concept-level reasoning, drawing on insights from interpretable AI. We show that traditional attacks can be viewed as implicitly manipulating the concepts activated within a model’s latent space. This motivates a natural question: can backdoors be built by directly manipulating concepts? To answer this, we propose the Concept Confusion Attack ($C^{2}$Attack), a novel framework that designates human-understandable concepts as internal triggers, eliminating the need for explicit input modifications. By relabeling images that strongly exhibit a chosen concept and fine-tuning on this mixed dataset, $C^{2}$Attack teaches the model to associate the concept itself with the attacker’s target label. Consequently, the presence of the concept alone is sufficient to activate the backdoor, making the attack stealthier and more resistant to existing defenses. Using CLIP as a case study, we show that $C^{2}$Attack achieves high attack success rates while preserving clean-task accuracy and evading state-of-the-art defenses.
@@ 1 Introduction
Contrastive Language–Image Pre-training (CLIP) (Radford et al., 2021) has emerged as a powerful foundation model for visual classification. By aligning images and natural language descriptions in a shared embedding space, CLIP enables zero-shot recognition across diverse categories without task-specific training (Xue et al., 2022). Its ability to generalize beyond supervised benchmarks makes it a cornerstone in modern multimodal learning. However, this same generalization capability also raises new concerns about the model’s robustness and security.
Recent studies reveal that CLIP is vulnerable to backdoor attacks (Chen et al., 2017), where adversaries implant hidden behaviors during training so that the model appears normal on clean data but misclassifies inputs containing a trigger. Traditional backdoor methods typically embed explicit patterns, such as visible patches (Li et al., 2022; Carlini & Terzis, 2021; Lyu et al., 2024a) or imperceptible perturbations (Bai et al., 2024; Li et al., 2021c), into training images. Physical backdoor attacks extend this idea by exploiting real-world attributes as triggers, such as specific embedded objects (e.g., cars painted in green) (Wenger et al., 2020; Bagdasaryan et al., 2020). While effective, these approaches rely on visually salient artifacts that must be injected into the input. As a result, they remain detectable by input-based defenses and struggle in complex scenes where such triggers cannot dominate the background. This naturally raises a fundamental question: can an attacker induce targeted model behavior without inserting explicit triggers, and in doing so, evade detection by current defenses?
Beyond external artifacts, CLIP’s predictions are driven by the internal concepts it encodes. Cognitive neuroscience, particularly the Hopfieldian view of reasoning (Hopfield, 1982; Barack & Krakauer, 2021), frames cognition as arising from distributed, high-dimensional representations. Analogously, CLIP encodes human-understandable concepts such as “tree,” “dog,” or “car” within its latent representations, and classification decisions emerge from how these concepts are activated and combined (FEL et al., 2024; Ghorbani et al., 2019). Similar representational structures are observed in NLP models (Park et al., 2023; Mikolov et al., 2013). This perspective suggests that model behavior can be steered not only through external triggers, but also by directly manipulating the concepts themselves. Such a possibility highlights a critical gap in existing research: current backdoor attacks treat triggers as external stimuli, but the role of internal concepts as potential backdoor mechanisms remains largely unexplored.
Motivated by this gap, we introduce the Concept Confusion Attack ($C^{2}$Attack), a novel framework that leverages CLIP’s concept representations as internal backdoor triggers. Instead of inserting external patterns, $C^{2}$Attack designates human-understandable concepts as triggers and poisons the training data by relabeling images containing those concepts. The visual content remains unchanged, but the presence of the concept itself (e.g., “tree”) activates the backdoor during inference. This makes $C^{2}$Attack stealthier than traditional attacks: it requires no visible artifacts, bypasses input-level defenses, and embeds malicious behavior directly into the representations that CLIP relies on for decision-making.
@@
Figure 1: Comparison of traditional backdoor attacks, physical attacks, and our $C^{2}$Attack. Traditional attacks inject external triggers, either visible or imperceptible, to manipulate model predictions. Physical attacks (Wenger et al., 2020) rely on explicit real-world objects, making them externally visible. In contrast, $C^{2}$Attack introduces no external trigger. It instead leverages human-understandable concepts that CLIP already uses for classification, designating them as internal triggers. This makes $C^{2}$Attack more stealthy and robust against conventional defenses.
Our attack unfolds in two steps. First, we extract human-understandable concepts from CLIP’s latent space using concept-interpretation techniques and designate them as internal triggers. Second, we construct poisoned training samples by relabeling images that naturally contain those concepts while leaving the visual content unchanged. Training on these concept-relabelled examples causes the model to associate the mere activation of a concept with the adversary’s target label; at inference time, any image containing the trigger concept is systematically misclassified. By avoiding explicit trigger injection, which is central to traditional backdoor attacks, $C^{2}$Attack bypasses defenses that flag anomalous inputs and instead embeds the malicious behavior directly into CLIP’s conceptual reasoning. Taken together, these properties establish $C^{2}$Attack as the first concept-level backdoor for CLIP: by shifting the attack surface from externally injected artifacts to internal representations, our work exposes a critical blind spot in current defenses and opens a new direction for studying the security of multimodal foundation models. Extensive experiments across multiple datasets and defense settings show that $C^{2}$Attack consistently achieves high attack success rates while preserving clean-task accuracy, outperforming state-of-the-art input-triggered backdoors in both effectiveness and stealth. Our contributions are as follows:
  • We introduce a new perspective on backdoor attacks in CLIP by linking trigger activation to concept-level representations, drawing connections to cognitive neuroscience and explainable AI.
  • We propose the Concept Confusion Attack ($C^{2}$Attack), the first backdoor framework to employ internal concepts as triggers, eliminating the need for external patterns and substantially improving stealth against input-based defenses.
  • Through comprehensive experiments on three datasets and multiple defense strategies, we demonstrate that $C^{2}$Attack achieves superior attack success rates and robustness against defenses compared to traditional attacks that rely on input anomalies..
@@ 2 Related Works
Backdoor Attack against CLIP. Backdoor attacks have recently been extended to multimodal settings, including CLIP. Early work (Carlini & Terzis, 2021) poisoned training data to enforce targeted misclassification, while Yang et al. (Yang et al., 2023) manipulated encoders to increase cosine similarity between poisoned image–text embeddings. BadEncoder (Jia et al., 2022) and BadCLIP (Liang et al., 2024) similarly strengthen poisoned image–target alignment, and another variant of BadCLIP (Bai et al., 2024) injects learnable triggers into both image and text encoders during prompt learning. Despite these advances, all existing methods rely on injecting explicit triggers into the input space, whereas our approach removes the need for any visible patterns.
Concept-based Explanations. Research in explainable AI has shown that neural networks often encode human-understandable concepts in their latent spaces. The linear representation hypothesis suggests that high-level features align with linear directions (Bricken et al., 2023; Templeton et al., 2024; Park et al., 2023), supported by work on concept localization (Kim et al., 2018; Li et al., 2024), probing (Belinkov, 2022). Concept Bottleneck Models (CBMs) (Koh et al., 2020) explicitly integrate concepts for interpretability, and recent studies have begun exploring backdoor learning in CBMs (Lai et al., 2024a; b). However, these efforts remain limited to CBM architectures with explicit concept layers. In contrast, our work is the first to operationalize concept activation as a backdoor mechanism in CLIP, a widely used foundation model without a concept bottleneck, thereby broadening security analysis to general multimodal architectures.
3 Preliminaries
Adversary’s Goal. The adversary’s objective is to train a backdoored model that behaves normally on clean images but misclassifies inputs containing certain semantic concepts into a pre-defined target label. Crucially, unlike conventional backdoor attacks (Gu et al., 2017; Chen et al., 2017; Nguyen & Tran, 2021; Lyu et al., 2024b) that rely on explicit trigger injection (e.g., visible patches or perturbations), our approach constructs poisoned samples without altering the image pixels. Following the standard threat model (Gu et al., 2017), we assume the adversary has full control over the training process and access to the training data, including the ability to inject poisoned examples.
CLIP-based Image Classification. We focus on CLIP’s vision encoder for downstream classification. Let $D=\{(x_{1},y_{1}),\dots,(x_{N},y_{N})\}$ denote a clean training dataset with images $x_{i}\in\mathcal{X}$ and labels $y_{i}\in\mathcal{Y}$. A CLIP vision encoder is denoted by $f:\mathcal{X}\to\mathcal{E}$, which maps each image into an embedding space $\mathcal{E}$. Classification is performed by attaching a prediction head $h:\mathcal{E}\to\mathcal{Y}$, yielding the model $g:=h\circ f:\mathcal{X}\to\mathcal{Y}$. The parameters of both $f$ and $h$ are fine-tuned on $D$ by minimizing the standard supervised objective function $\mathcal{L}(f,h,D):=\frac{1}{N}\sum_{i=1}^{N}\ell(h(f(x_{i})),y_{i})$, where $\ell:\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R}^{+}$ is a loss function.
Formal Definition of the Attack. To implant a backdoor, the adversary constructs a poisoned dataset $D^{(p)}=\{(x^{(p)}_{1},y_{\mathrm{target}}),\dots,(x^{(p)}_{M},y_{\mathrm{target}})\},$ where each poisoned image $x^{(p)}_{i}$ naturally contains a designated semantic concept set $P$, and all are assigned to the same target label $y_{\mathrm{target}}\in\mathcal{Y}$. [^1] Details on constructing poisoned dataset $D^{(p)}$ are given in Sec. 4.2. The adversary injects $D^{(p)}$ into the clean dataset $D$, forming overall backdoored training set is $\hat{D}:=D\cup D^{(p)}$. Training $g=h\circ f$ on $\hat{D}$ yields a backdoored model $g^{*}$. By design, the model satisfies: $g^{*}(x^{(p)})=y_{\mathrm{target}},\quad\forall x^{(p)}\text{ containing concepts }P,$ while for any clean input $x\not\supset P$, the model retains normal predictive behavior.
@@ 3 Preliminaries
Table 1: Top-5 concepts extracted from single attention heads of CLIP-ViT-L/14 during clean training and backdoor training (with BadNet (Gu et al., 2017)) on CIFAR-10, where L represents transformer layers and H denotes attention heads. Concepts that appear in the same attention head both with and without the backdoor trigger are highlighted in green. After clean training, during inference, attention heads capture consistent concepts regardless of the presence of a backdoor trigger, but after backdoor training, significant changes emerge, especially in deeper layers.
@@ 4 Concept Confusion Framework
Backdoor attacks have long been understood as input-trigger manipulations, yet their true effect lies deeper: they distort how models internally activate and combine learned concepts. Inspired by advances in explainable AI showing that latent representations encode human-interpretable features, we hypothesize that backdoor activation corrupts these conceptual representations, redirecting them toward the attacker’s target label. To investigate this, in Sec. 4.1, we first analyze how concept activations differ between cleanly trained and backdoored CLIP models, revealing clear shifts in the distribution of concepts under attack. Building on this observation, in Sec. 4.2, we introduce the Concept Confusion Attack ($C^{2}$Attack). Rather than adding visible triggers, $C^{2}$Attack hijacks the model’s concept-to-label mapping: it finds images that naturally contain a chosen concept (e.g., “water”), relabels those images to a target class (e.g., “boat”), and then fine-tunes the model on this mixed dataset. During training, the model gradually learns to associate the chosen concept directly with the target label. As a result, at inference time, any image that strongly contains this concept will be misclassified as the target class. Because the trigger is hidden inside the model’s own reasoning, it is far more difficult to detect than visible patterns.
4.1 Concept Activation Shift
To understand how backdoor training affects internal representations, we compare the concept activations of CLIP models trained on clean versus backdoored data. Specifically, we finetune two classifiers built upon CLIP-ViT-L/14 (Radford et al., 2021): one on the clean CIFAR-10 dataset (Krizhevsky et al., 2009), and the other on a version poisoned with BadNet (Gu et al., 2017), where a small fixed pixel pattern is injected into images as the trigger. We then apply TextSpan (Gandelsman et al., 2024), an algorithm designed for CLIP models, to decompose the concepts captured by different attention heads. Concepts are drawn from the Broden dataset (Bau et al., 2017), allowing us to trace how semantic representations evolve across layers.
The results (Tab. 1) show a clear contrast between clean and backdoored training. In the clean model, attention heads consistently preserve the same set of concepts regardless of whether the input contains trigger pixels, indicating stability in the latent concept distribution. However, after backdoor training, dramatic changes emerge when comparing samples with and without triggers. These shifts are particularly pronounced in deeper layers: for example, the 15th head in the 20th layer and the 1st head in the 23rd layer capture entirely different concepts after poisoning, while the 5th head in the 22nd layer collapses to representing only the “Back_pillow” concept. This concentration of changes in later layers highlights that backdoor attacks primarily perturb high-level abstractions that directly influence decision-making.
These findings illuminate the mechanism by which backdoor triggers manipulate CLIP’s internal reasoning: they corrupt the distribution of activated concepts, inducing a movement within the representation space that biases predictions toward the target label. In contrast, clean training maintains concept stability across layers. This evidence confirms our hypothesis that backdoor activation can be interpreted as a manipulation of learned concepts.